### v1

# Structures resistant to Manipulation by all Wavefronts

Abstract
Using light to manipulate small particles is a tool with many practical applications throughout biophysics and nanotechnology. These tools have seen a significant increase in performance by utilizing shaped wavefronts, most commonly created with spatial light modulators. Wavefront shaping has also enabled the manipulation of seemingly arbitrary objects, which was impossible with conventional beams. In contrast, we show here the existence of a wide variety of objects that cannot be manipulated as desired, even with the optimal wavefront shaping protocol. The counterintuitive shapes of these objects are found using inverse design. Specifically, we show that the maximal pulling force is reduced by up to three orders of magnitude, and the maximal trapping stiffness is reduced by up to nearly two orders of magnitude. Our findings could prove useful in the development of micromachines that require a predictable mechanical reaction to an arbitrary wave.
Introduction
Since the pioneering work of Ashkin, who first demonstrated that tightly focused light can stably trap microscopic particles [ashkin_observation_1986], optical micromanipulation has developed into a versatile tool across physics, biology, and soft matter [padgett_tweezers_2011, dholakia_shaping_2011, volpe_roadmap_2023]. For example, optical trapping has been used to elucidate the biomechanics of biological processes [block_bead_1990, heller_optical_2014], control micromachines [grier_revolution_2003, butaite_indirect_2019], experimentally validate the connection between entropy and information [berut_experimental_2012, saha_maximizing_2021], and aid in cooling nanoparticles to their motional ground state [delic_cooling_2020, tebbenjohanns_quantum_2021]. The same optical fields can also transfer angular momentum [friese_optical_1998] and even generate pulling forces [novitsky_single_2011, brzobohaty_experimental_2013]. Related capabilities also exist in acoustics, including acoustic tweezers for manipulating single microparticles, cells, and entire organisms [ding_-chip_2012] and acoustic tractor beams [demore_acoustic_2014].
The ability to shape the wavefront of light [mosk_controlling_2012, rotter_light_2017, forbes_structured_2021, bertolotti_imaging_2022] beyond the conventionally used Gaussian beams has advanced micromanipulation capabilities [marzo_holographic_2015, otte_optical_2020]. For example, wavefront shaping enables trapping behind scattering layers [cizmar_situ_2010], stiffer traps [taylor_enhanced_2015, butaite_photon-efficient_2024], controlled transfer of angular momentum [stilgoe_controlled_2022], trapping at multiple spots [garces-chavez_simultaneous_2002], and trapping and manipulating rod-shaped bacteria [lenton_orientation_2020]. Alternatively, the near-field around the particle can be tailored by inverse-designing nanocavities and plasmonic apertures [martinez_de_aguirre_jokisch_omnidirectional_2024, nelson_inverse_2024]. Acoustic transducer arrays provide an analogous form of control and can, for example, drive objects even through disordered and dynamic media [orazbayev_wave-momentum_2024]. A complementary route for improving micromanipulation is inverse-designing particles for fixed wavefront illumination [phillips_shape-induced_2014, stein_shaping_2022, lee_inverse_2025]. Bounds on optimal forces, torques, and illumination fields [liu_optimal_2019, kuang_maximal_2020] solidified the understanding that shaped waves have a large influence for manipulating objects. These successes raise a fundamental question: are there targets for which mechanical control remains fundamentally limited even with optimal wavefront shaping?
In this paper, we identify targets for which even the optimal incident wavefront cannot efficiently realize certain micromanipulation tasks. We inverse-design structures whose mechanical response is predictable for an arbitrary wavefront (within fixed illumination constraints). As one example, we numerically show that the maximum achievable pulling force (“optimal tractor beam”) on these inverse-designed targets can be suppressed by up to three orders of magnitude relative to a generic structure with a uniform refractive index distribution. Remarkably, the ability to design such structures persists across broad parameter ranges, indicating that wavefront shaping does not guarantee arbitrary mechanical control. Our results not only show that it is possible to find structures that are predictably manipulated by an arbitrary wave, enabling reliable micromanipulation with dynamically distorted wavefronts. They also demonstrate the existence of fundamental limits to improving micromanipulation with wavefront shaping.
Figure 1: Reduction of optimal tractor beam strength with inverse design. Intensity distribution resulting from the optimal tractor beam for a conventional structure (a) with a uniform refractive index distribution ($n=3$) and for an inverse-designed structure (b) to minimize the optimal pulling force. Here, the maximal pulling force ($F$) on the inverse-designed structure is reduced by nearly two orders of magnitude compared to the conventional structure. We restrict the aperture of the incoming wave to one-quarter of the circular boundary. Insets display the refractive index distributions of both structures and the modal distributions of the optimal tractor beam wavefronts. The gray square indicates the extent of the design region.
Figure 2: Broad parameter range for tractor beam and trapping resistant particles. Plots show by how much the optimal micromanipulation has been reduced for an inverse-designed target compared to a conventional structure with a uniform refractive index. The top row and bottom give the reduction factor for the optimal tractor beam (i.e., the largest eigenvalue of $Q_{x}$) and optimal trapping stiffness (i.e., the largest eigenvalue of $W_{x}$), respectively. The reduction factor is shown for different distances (a, e), widths of the design region (b, f), maximal refractive index of the structure (c, g), and wavelength (d, h). Insets show examples of inverse-designed structures that minimize this ratio. The black dashed line indicates the parameter value used in the other scans.
Inverse design of manipulation-resistant structures
To inverse-design structures where the optimal wavefront exhibits a negligible influence on the target, we first need the ability to compute the optimal wavefront. For the first example of targets designed for greatly reduced optimal pulling forces, we employ the generalized Wigner-Smith operator $Q_{x}$, which was introduced for optimal focusing on targets in disordered media [ambichl_focusing_2017] and has in the meantime been applied to a wide range of manipulation tasks [horodynski_optimal_2020, horodynski_tractor_2023, butaite_photon-efficient_2024, hupfl_optimal_2023]. The optimal tractor beam wavefront is found by solving the eigenvalue problem of $Q_{x}$ (which is Hermitian in the absence of loss and gain in the system) for the largest eigenvalue, which reads
$Q_{x}\vec{u}_{x}=(K_{x}^{\text{in}}-S^{\dagger}K_{x}^{\text{out}}S)\vec{u}_{x}=q_{x}\vec{u}_{x},$ (1)
where $S$ is the target’s scattering matrix, $K_{x}^{\text{in}}$ ($K_{x}^{\text{out}}$) describes the incoming (outgoing) momentum of an arbitrary wavefront, $\vec{u}_{x}$ is a vector of modal amplitudes, and $q_{x}$ is the eigenvalue. From this equation, we can interpret $Q_{x}$ as describing the difference between the wave’s incoming and outgoing momentum, which is transferred (by conservation of momentum) to the target. We restrict our analysis to two spatial dimensions, where the scattering of waves is governed by the Helmholtz equation $\Delta+k^{2}n^{2}(\textbf{r})]\psi(\textbf{r})=0$, with $\Delta$ the Laplacian, $k=|\textbf{k}|$ the wavevector, $\psi(\textbf{r})$ the scalar wave, and $n(\textbf{r})$ is the spatially varying refractive index (which is continually adjusted during the inverse design process). We solve the two-dimensional Helmholtz equation numerically using our publicly available code [horodynski_open_2023] based on an open-source finite-element method (NGSolve) [schoberl_netgen_1997, schoberl_c11_2014]. We note that our approach can be straightforwardly extended to three dimensions and two polarizations (though at a higher computational cost).
Our ultimate goal is to find structures for which the largest eigenvalue of $Q_{x}$, $q_{x}^{\text{max}}$ is vanishingly small. Since this eigenvalue gives — for all possible incoming wavefronts — the largest pulling force (i.e., the strongest tractor beam), $q_{x}^{\text{max}}\sim 0$ implies that it is impossible to pull the target with any wave. We achieve this goal by employing inverse design, in which we continuously adjust the refractive index distribution of the structure (parameterized by a square grid of pixels with individually tunable refractive indices between a minimum and maximum value) guided by a cost function (which is just $q_{x}^{\text{max}}$). To use state-of-the-art gradient-based inverse design algorithms, we require a computationally efficient way of calculating the gradient of the cost function with respect to the refractive index, $n_{i}$, of each design pixel, i.e., $\partial_{n_{i}}q_{x}^{\text{max}}=\left(\vec{u}_{x}^{\text{max}}\right)^{\dagger}\partial_{n_{i}}Q_{x}\vec{u}_{x}^{\text{max}}$. From Eq. (1), we see that calculating $\partial_{n_{i}}Q_{x}$ amounts to the computation of $\partial_{n_{i}}S$, since $K_{x}$ is independent of the refractive index distribution. We efficiently calculate the sensitivity of the scattering matrix with respect to the refractive indices of the design pixels by using a method based on an alternative form of the generalized Wigner-Smith operator naturally suited to the task at hand [horodynski_anti-reflection_2022]. We refer to the Supplementary Information for a detailed description of the gradient computation and inverse design algorithm.
In Fig. 1, we show the resulting structure of one such design run and compare its performance to that of a conventional structure. More specifically, we place a square design region ($0.4\lambda\times 0.4\lambda$) with a grid of $60\times 60$ design pixels in the middle of a circular scattering region. We can inject wavefronts into the scattering region with an arbitrary amplitude and phase distribution in a $\pi/2$ solid angle. These wavefronts always have unit norm for comparable results, i.e., $\vec{u}^{\dagger}\vec{u}=1$. Our simulations reveal an optimal pulling force on the conventional structure (a square with a uniform refractive index of $n=3$) of $F_{x}\sim 1.2$ (a.u.). In contrast, the optimal pulling force on the inverse-designed structure is smaller by nearly two orders of magnitude: $F_{x}\sim 0.016$ (a.u.). This drop in pulling force is achieved despite comparable intensity to the baseline case in the design region. We note that the wave’s ability to apply a pushing force is (in principle) unaffected by our design procedure. In this example, the maximal pushing force on the inverse-designed [conventional] target is $F\sim-89.4$ (a.u.) $F\sim-84.2$ (a.u.)].
Tractor beam resistant structures
Having presented one example of an inverse-designed structure where the optimal pulling force is reduced by two orders of magnitude, we now turn our attention to the question of whether such structures can be found generically. To this end, we vary various parameters of the studied system and investigate how much the optimal pulling force is reduced compared to a corresponding conventional structure. Concretely, we sweep the distance of the structure to the wave’s source, the side length of the square design region, the maximum refractive index, and the wavelength. We note that each structure is designed for one particular wavelength, i.e., for monochromatic illumination, and that changing the distance of the structure to the source also indicates the influence of varying the numerical aperture.
Our results are summarized in Fig. 2a-d and clearly show that structures (strongly) resistant to the optimal pulling force occur over a wide range of parameters. This points to the conclusion that we have uncovered a large class of structures and that wavefront shaping does not generically guarantee arbitrary micromanipulation capabilities. Our data also clearly indicates that reducing the effective size of the design region (either by changing the size itself, or indirectly by increasing the wavelength or decreasing the maximal refractive index) results in a decreasing ability of the inverse-design algorithm to identify structures that perform markedly better than the conventional structures. Intuitively, this observation is explained by the transition to purely dipole-scattering for targets that are much smaller than the wavelength. In this case, the patterning is not resolved by the wave, and thus inverse design loses its influence over the optimal pulling force. These observations stay consistent if we restart the inverse design processes with a different set of randomly chosen refractive indices in the design region. The most noteworthy deviation from this general trend was for a particular size ($0.53\lambda$) of the design region, for which the performance was markedly worse than for comparable sizes. This spike appeared consistently, even after randomizing the initial guess of the inverse design process three times per data point and running the inverse design process for more iterations. A possible explanation is a resonance effect, as the size in question is close to $\lambda/2$.
[…]
Structures resistant to optical trapping
We now show that the inverse design of structures resistant to manipulation also extends to the optical trapping of particles. Accordingly, we use an operator derived from the generalized Wigner-Smith operator in Eq. (1) that allows us to compute the wavefront with the optimal trapping stiffness. The trapping stiffness (in $x$-direction) $\kappa_{x}=-\partial_{x}F_{x}$ quantifies how tightly a particle is trapped by computing how quickly a restoring force builds when the particle is displaced from its equilibrium position. The corresponding operator is given by [butaite_photon-efficient_2024]
$W_{x}=-\partial_{x}Q_{x}=i\left(S^{\dagger}K_{x}^{\mathrm{out}}SK_{x}^{\mathrm{in}}-K_{x}^{\mathrm{in}}S^{\dagger}K_{x}^{\mathrm{out}}S\right),$ (2)
where we used $\partial_{x}S=i(SK_{x}^{\mathrm{in}}-K_{x}^{\mathrm{out}}S)$ [ambichl_focusing_2017] to arrive at $W_{x}$. Since the only part of $W_{x}$ that depends on the refractive index of the design pixels is the scattering matrix $S$, we employ the same inverse design strategy as for the tractor beam case. Here, we use a circular design region to connect our results to optical trapping experiments, which usually employ silica nanospheres as targets. This also means that the conventional structure used as the baseline is circular with a uniform refractive index distribution.
For structures designed to resist an optical trapping beam, we find a wide range of parameters in which the performance of the optimal trapping beam degrades by up to two orders of magnitude relative to a conventional structure (see Fig. 2e-h). The broad trend of effectively larger structures performing better at resisting a trapping beam is similar to the case of structures designed to withstand the optimal pulling force. Interestingly, with increasing distance to the wave’s source, the performance of trapping-resistant targets improves, while the performance of tractor beam-resistant targets stays unchanged. In both micromanipulation examples considered here, the structures found with our inverse-design process consist mostly of two materials, as we employed hyperbolic tangent regularization to guide the optimizer to solutions exhibiting this property. The structures also exhibit comparably smooth boundaries without nudging them towards this property in the inverse-design process.
However, in optical trapping, it is usually not sufficient to only target improved trapping stiffness with wavefront shaping [volpe_roadmap_2023]. Rather, great care has to be taken to ensure that a trap is not only stiff, but also stable, i.e., no force is applied at zero displacement ($x=0$). In our case, this means that the trapping beam should not only maximize $\vec{u}^{\dagger}W_{x}\vec{u}$, but also ensure that $\vec{u}^{\dagger}Q_{x}\vec{u}=0$ at the equilibrium position, where $\vec{u}$ is the optimal wavefront. We find wavestates that possess these two properties by numerical optimization (see SI for a detailed description of the procedure), as no simple eigenproblem is available to compute the corresponding wavefront [butaite_photon-efficient_2024].
We analyze the response of conventional and inverse-designed structures to wavefronts that maximize trapping stiffness while keeping the trap stable in Fig. 3. Our results reveal that changing the wavefront from one that just maximizes trapping stiffness to one that also results in stable trapping does not adversely affect the purpose of the inverse-designed structure. I.e., the optimal trapping stiffness is still reduced by approximately an order of magnitude, and the potential well created by the optical trap is much shallower for the inverse-designed structure.
Outlook and discussion
To summarize, we have inverse-designed structures that are resistant to pulling forces and optical trapping beams, even when the optimal wavefront for the task at hand is used. We showed the viability of our concept via numerical simulations, reducing manipulation capabilities on inverse-designed structures by up to three orders of magnitude compared to conventional structures. The properties of targets we found potentially enable the application of our inverse-design pipeline to the fabrication of structures suited both for optical [lee_inverse_2025], and acoustic [stein_shaping_2022] micromanipulation. Although the results presented here are for a two-dimensional system, the extension to three-dimensional vector waves should be possible.
Our approach offers an interesting perspective on optically driven micromachines [zhang_fabrication_2022], when only specific degrees of freedom need to be addressed. It potentially enables the operation of such micromachines even when they are embedded in a disordered, changing environment, i.e., when the illuminating wavefront is constantly evolving. Manipulation-resistant structures could also be used to reduce back-action noise in optical levers [hao_back_2024].
Acknowledgements
This research was funded in part (M.H.) by the Austrian Science Fund (FWF) [10.55776/J4729]. This material is based upon work also supported in part by the U. S. Army Research Office through the Institute for Soldier Nanotechnologies at MIT, under Collaborative Agreement Number W911NF-23-2-0121. The authors acknowledge the MIT SuperCloud and Lincoln Laboratory Supercomputing Center for providing high-performance computing resources that have contributed to the research results reported within this paper.